Marcelino Co.s March 31 inventory of raw materials is $86,000. Raw materials purchases in April are $560,000, and factory paa. Materials purchases (on credit). b. Direct materials used in production. c. Direct labor paid and assigned to Work in Proc3. Prepare a schedule of cost of goods manufactured. MARCELINO COMPANY Schedule of Cost of Goods Manufactured For Month Ended4.1 Compute gross profit for April. Gross Profit
Answer #1

Solution:

2.) Journal entries

Transaction General journal Debit ($) Credit($)
a.) Raw materials inventory 560,000
Accounts payable 560,000
b.) Work in process inventory (132,000+215,000+110,000) 457,000
Raw material inventory 457,000
c.) Work in process inventory (101,000 + 150,000 + 104,000) 355,000
Cash 355,000
d.) Factory overhead 23,000
Cash 23,000
e.) Work in process inventory (355000 of 50%) 177,500
Factory overhead 177,500
f(1) Factory overhead 51,000
Raw materials inventory 51,000
f(2) Factory overhead 24,000
Cash 24,000
f(3) Factory overhead 53,000
Accumulated depreciation -factory equipment 53,000
f(4) Factory overhead 33,000
Cash 33,000
g.) Finished good inventory 845,000
Work in process inventory 845,000
h.) Cost of goods sold 343,500
Finished goods inventory 343,500
I) Cash 630,000
Sales 630,000
j) Cost of goods sold 6,500
Factory overhead 6,500

g.) beginning work in process +Direct material + direct labor +Applied overhead)

For job 306 =(27,000 +22,000+11,000 +132,000+101,000+ 50,500) =343,500 (goods sold)

For job 307 =(36,000+17,000 +8,500 +215,000 +150,000 +75,000) =501,500

Finished goods inventory =343,500+501,500 =845,000

j. Factory overhead

Overhead applied to jobs $177,500
Overhead incurred:
Indirect material $51,000
Indirect labor $23,000
Factory rent $33,000
Factory utilities $24,000

Factory depreciation on equipment =53,000

51,000,+23,000+33,000+24,000+53,000 =$184,000

Under applied overhead =184,000-177,500 =6,500

3.) Marcelino company: Schedule of cost of goods manufactured

For month ended April

Direct material used $457,000
Direct labor used $355,000
Factory overhead applied $177,500
Total manufacturing cost $989,500
Add: work in process March 31 $121,500
Total cost of work in process $1,111,000
Less:work in process April 30(110,000+104,000+52,000) 104000 of 50%=52,000 $266,000
Cost of goods manufactured $845,000

4.) Gross profit =$280,000

Sales $630,000
Less: cost of goods sold (343500+6500) $350,000
Gross profit $280,000

*sorry for inconvenience of( j.)